A 12-oz serving of mountain Dew contains 55 mg of caffeine how much caffeine is in an 18-oz serving of mountain dew?
Found 3 solutions by Alan3354, stanbon, amalm06:
Answer by Alan3354(69443)

18 oz is 1.5 times as much.
--> 1.5 times the caffeine

Answer by stanbon(75887)

A 12-oz serving of mountain Dew contains 55 mg of caffeine how much caffeine is in an 18-oz serving of mountain dew?
--------
Use a proportion::
x/18 = 55/12
x = (18/12)55
x = (3/2)55
x = 82.5 mg
----------------
Cheers,
Stan H.
----------------

Answer by amalm06(224)

(55/12)(18)=82.5 mg (Answer)